### v2

@@ 8 Generalized network utilities
We claim that $\delta^{\prime}>1$. Recall that fitting pairs are required to satisfy (9). The sign of $\gamma_{e}\pi^{\prime}_{w}-\pi^{\prime}_{v}$ is the same as the sign of $\gamma_{e}\pi_{w}-\pi_{v}$ for all arcs $e=(v,w)$ with both endpoints inside $T$ or outside $T$. There is no arc $e=(v,w)$ entering or leaving $T$ with $0<f^{\prime}_{e}<u_{e}$, since $e,\overleftarrow{e}\in E^{\textrm{aux}}_{f^{\prime},\pi}$ for all such arcs. For every arc $e=(v,w)$ entering $T$ with $f^{\prime}_{e}=0$, we must have $\gamma_{e}\pi_{w}-\pi_{v}<0$ so that $e\notin E^{\textrm{aux}}_{f^{\prime},\pi}$. Similarly, for every arc $e=(v,w)$ leaving $T$ with $f^{\prime}_{e}=u_{e}$, we must have $\gamma_{e}\pi_{w}-\pi_{v}>0$. Hence, we can always choose $\delta^{\prime}>1$. If $\delta^{\prime}=\alpha^{\prime}$, then $\pi^{\prime}$, $x^{\prime}$, $\mathbf{q}$ and $b^{\prime}$ satisfy (10), since the right hand side for all $j\in\operatorname{supp}(x^{\prime})$ has increased by a factor $\alpha^{\prime}$, and therefore equality holds for all these indices. We can therefore conclude $x^{\prime}\in\mathcal{GD}^{u}(b,\mathbf{q})$ and terminate by outputting $x^{\prime}$.
@@
In case $\delta^{\prime}<\alpha^{\prime}$, then (10) holds with strict inequality for all $j$ for $\pi^{\prime}$, $x^{\prime}$, $\mathbf{q}$, and $b^{\prime}$. Moreover, the ratio of the right hand side and the left hand side is the same for all $j\in\operatorname{supp}(x^{\prime})\cap\operatorname{supp}(\mathbf{q})$, namely, $\alpha^{\prime}/\delta^{\prime}<\alpha$. We now restart the above procedure with $(f^{\prime},\pi^{\prime})$ and $x^{\prime}$ in place of $(f,\pi)$ and $x$.
@@
We claim that the algorithm terminates in a finite number of iterations. This holds because the linear program Pullback-Flow only depends on the sign pattern of $\gamma_{e}\pi_{w}-\pi_{v}$ for the arcs $e=(w,v)$ and on $\operatorname{supp}(x)$. There are $3^{|E|}$ possible sign patterns; the sest $\operatorname{supp}(x)$ is monotone decreasing and therefore there are at most $|{\mathcal{M}}|$ different possibilities. The same configuration of arc sign patterns and $\operatorname{supp}(x)$ cannot occur twice, because $u(x^{\prime})=\operatorname{net}_{f^{\prime}}(t)$ is strictly decreasing in every iteration of the algorithm. Hence, after a finite number of iterations we arrive at $\pi^{\prime}$ and $x^{\prime}$ satisfying (10) for $b^{\prime}$. Since $x^{\prime}$ is decreasing throughout, we have $x^{\prime}\leq x$ for the input supply vector $x$, as required. ∎
Input: $\mathbf{q}\in\operatorname{pdom}(u)$, $b,b^{\prime}>0$, $x\in\mathcal{GD}^{u}(b,\mathbf{q})$.
Output: $0\leq x^{\prime}\leq x$ such that $x^{\prime}\in\mathcal{GD}^{u}(b^{\prime},\mathbf{q})$.
@@
1 Find a fitting pair $(f,\pi)$ with respect to $x$ such that $u(x)=\operatorname{net}_{t}(x)$ and (10) holds;
2 if $u(x)=0$ or $\mathbf{q}=\mathbf{0}$ then return $x$;
3 $\bar{x}\leftarrow x$ ; $\alpha\leftarrow b/b^{\prime}$ ;
@@
5       $(x^{\prime},f^{\prime})\leftarrow\textnormal{{Pullback-Flow}}(\bar{x},f,\pi)$ ;
6       $\alpha^{\prime}\leftarrow\alpha\frac{u(x^{\prime})}{u(x)}$ ;
@@ 8             $\lambda\leftarrow\frac{\alpha-1}{\alpha-\alpha^{\prime}}$ ;
9             $x^{\prime\prime}\leftarrow\frac{\alpha-1}{\alpha-\alpha^{\prime}}x^{\prime}+\frac{1-\alpha^{\prime}}{\alpha-\alpha^{\prime}}\bar{x}$ ;
10             return $x^{\prime\prime}$
@@
12             $T\leftarrow\left\{v\in V\,:\,v\text{ can reach }\operatorname{supp}(x^{\prime})\text{ in }(G,E^{\textrm{aux}}_{f^{\prime},\pi})\right\}$  ;
13             $\delta\leftarrow\min\left\{\alpha,\min\left\{\frac{\gamma_{e}\pi_{w}}{\pi_{v}}\,:\,e=(v,w)\in\partial^{\mathrm{in}}(T)\right\},\min\left\{\frac{\pi_{v}}{\gamma_{e}\pi_{w}}\,:\,e=(v,w)\in\partial^{\mathrm{out}}(T)\right\}\right\}$;
14             if $\delta^{\prime}=\alpha^{\prime}$ then return $x^{\prime}$;
@@
16             $\alpha\leftarrow\alpha^{\prime}/\delta^{\prime}$ ; $\bar{x}\leftarrow x^{\prime}$ ;
@@
Lemma 8.8.
Let $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ be a generalized network utility function and $\mathbf{q},\mathbf{q}^{\prime}\in\operatorname{pdom}(u)$ such that $0\leq\mathbf{q}^{\prime}\leq\mathbf{q}$, and let $x\in\mathcal{GD}^{u}(b,\mathbf{q})$. Then, there exists $x^{\prime}\in\mathcal{GD}^{u}(b,\mathbf{q}^{\prime})$ such that $x^{\prime}_{j}\leq x_{j}$ whenever $q^{\prime}_{j}=q_{j}$.
@@
Clearly, it suffices to show that the statement holds when $q^{\prime}_{k}<q_{k}$ for a single good $k\in{\mathcal{M}}$. Again, we start from fitting pair $(f,\pi)$ with respect to $x$ as in Lemma 8.6. Similarly to Lemma 8.7, we may assume that $u(x)>0$ and $\mathbf{q}\neq\mathbf{0}$. Let us now replace $\mathbf{q}$ by $\mathbf{q}^{\prime}$. The condition (10) can only be violated for $j=k$. If it is not violated, then we may return $x^{\prime}=x$; this is always the case when $\pi_{k}=0$. For the rest of the proof, assume that the (10) is violated for $k$ with the cost $q^{\prime}_{k}$. Note that $\pi_{j}=\beta q^{\prime}_{j}$ for every $j\neq k$, where $\beta\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}u(x)/b$; we now have $\pi_{k}>\beta q^{\prime}_{k}\geq 0$. This value of $\beta$ will be fixed throughout the argument.
@@
We can again argue that this algorithm terminates in a finite number of iterations. Note that $\operatorname{supp}(x^{\prime})$ is monotone decreasing throughout, and $\pi^{\prime}_{j}=\beta q^{\prime}_{j}$ is maintained for all $j\in\operatorname{supp}(x^{\prime})\setminus\{k\}$. For goods with $x^{\prime}_{j}=0$, $\pi^{\prime}_{j}$ may decrease, and therefore $\pi^{\prime}_{j}\leq\beta q_{j}$ for these goods. At termination, we have $\pi^{\prime}_{j}=\beta q^{\prime}_{j}$ for all $j\in\operatorname{supp}(x^{\prime})$ — including $j=k$ — and $\pi^{\prime}_{j}\leq\beta q^{\prime}_{j}$ whenever $x^{\prime}_{j}=0$. Moreover, note that $x^{\prime}_{j}\leq x_{j}$ for all $j\neq k$.
Throughout the algorithm, $u(x^{\prime})$ is non-decreasing. Hence, $u(x^{\prime})/b\geq\beta$ holds. By setting $\hat{b}>b$ such that $u(x^{\prime})/\hat{b}=\beta$, (10) holds for $x^{\prime}$, $\pi^{\prime}$, $q^{\prime}$ and $\hat{b}$. Hence, $x^{\prime}\in\mathcal{GD}_{u}(\hat{b},\mathbf{q}^{\prime})$. Applying now Lemma 8.7 for $x^{\prime}$, the prices $\mathbf{q}^{\prime}$ and budgets $\hat{b}\geq b$, we can find $x^{\prime\prime}\in\mathcal{GD}_{u}(b,\mathbf{q}^{\prime})$ such that $x^{\prime\prime}\leq x^{\prime}$. Note that $x^{\prime\prime}_{j}\leq x^{\prime}_{j}\leq x_{j}$ for all $j\neq k$, and therefore $x^{\prime\prime}$ satisfies the requirements of the Lemma.
@@
This concludes the proof of Theorem 1.5, namely, that all generalized network utilities are Gale-substitutes. Nevertheless, they are not $\Sigma$-Gale-substitutes, and so Theorem 1.3 is only applicable in the case when the generalized network utilities are non-satiable. Here, we also show Theorem 1.6, namely, that the desired utility bound holds for all generalized network utilities if we consider thrifty competitive equilibria.
See 1.6
Proof sketch.
We only need to slightly modify the argument in Section 6. Property ($\Sigma$) is only used in the proof of Lemma 3.2. We show that for generalized network utilities, if $x\in{\mathcal{D}}^{u}(\mathbf{p},b)$ for $\mathbf{p}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ and further $x$ minimizes $\left\langle\mathbf{p},x\right\rangle$ in ${\mathcal{D}}^{u}(\mathbf{p},b)$, then there exists $y\in\mathcal{GD}^{u}(\mathbf{p},b)$ such that $y\leq x$. From this property, we can derive Lemma 6.1 for thrifty competitive equilibria; the rest of the proof remains identical to the proof of Theorem 1.3.
Consider now (6) in the proof of Lemma 3.2. If $\lambda>0$, the same argument is applicable. Consider now the case $\lambda=0$. This means that the constraint $\left\langle\mathbf{p},x\right\rangle\leq b$ is not active in the program defining ${\mathcal{D}}^{u}(\mathbf{p},b)$; that is, $\max_{z\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}\{u(z)\,:\,\left\langle\mathbf{p},z\right\rangle\leq b\}=\max_{z\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}u(z)=v^{\star}$. In particular, $u(x)=v^{\star}$. The budget minimization property means that $x$ is an optimal solution to the program
$\min\leavevmode\nobreak$ $\mathbf{p}^{\top}x$ (11)
$\operatorname{net}_{f}(t)$ $\geq v^{\star}$
$\operatorname{net}_{f}(v)$ $\geq-x_{v}\quad\forall v\in S$
$\operatorname{net}_{f}(v)$ $\geq 0\quad\forall v\in V\setminus(S\cup\{t\})$
00 0 $\leq f\leq h\,.$
From here, using LP duality and Lemma 8.1, we can see that there exists a supergradient $\mathbf{g}\in\partial^{\wedge}u(x)$ and $\mu\in\mathbb{R}_{\geq 0}$ such that
@@
If $\mu>0$, then we can use the same argument for $\lambda=1/\mu$. If $\mu=0$, given that $\mu$ acts as the dual variable of $\operatorname{net}_{f}(t)\geq v^{\star}$, which is no longer active, the optimal value of (11) becomes 00 since $\mathbf{0}$ is a feasible allocation of $u(\cdot)$ according to our definition. This implies the existence of a zero-cost optimal allocation $x$: $u(x)=v^{*}$ and $\mathbf{p}^{\top}x=0$, which indicates that $x$ belongs to both ${\mathcal{D}}^{u}(\mathbf{p},b)$ and $\mathcal{GD}^{u}(\mathbf{p},b)$. Thus, the result follows. ∎
9 Discussion
In this paper, we compared competitive equilibrium and Nash welfare maximizing solutions in the Fisher market setting. We identified Gale-substitutes as a broad family of utility functions where the Nash welfare uniformly guarantees every agent at least half of their maximum utility at any competitive equilibrium. We also introduced generalized network utilities. This is a very general class with a rich combinatorial structure and may be a relevant example in various market and fair division settings.
Gale-substitutes may be a class of concave functions worth investigating on its own. In the context of indivisible goods, the analogous class of gross substitute utilities, introduced by Keslo and Crawford (1982), turned out to be the same as the important class of $M^{\natural}$-concave utilities in discrete convex analysis Fujishige and Yang (2003). In particular, one may wonder whether the Gale-substitutes property implies submodularity, that is, $u(x)+u(y)\geq u(x\vee y)+u(x\wedge y)$ for any $x,y\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$, where $x\vee y$ and $x\wedge y$ denote the pointwise maximum and minimum of the two vectors, respectively. This property trivially holds for separable utilities and can also be shown for generalized network utilities.
We note that one can construct an example of a concave quadratic function that is submodular but not Gale-substitutes. On the other hand, it is unclear whether Theorem 1.6 would remain true for submodular utilities instead of Gale-substitutes. One would also be interested in whether Theorem 1.6 holds for matching utilities, which, like submodular utilities, are not necessarily Gale-substitutesl(see Example A.11). Another natural question related to the well-studied notion of weak gross substitutes (WGS) utilities: Gale-substitutes includes utility functions that are not WGS, such as concave separable; but it is not clear whether all WGS utilities are also Gale-substitutes.
Gale-substitutes utilities may also have potential algorithmic applications. The WGS property leads to natural tâtonnement and auction mechanisms for finding competitive equilibria, see, e.g., Arrow et al. (1959), Arrow and Hurwicz (1960), Bei et al. (2019a), Codenotti et al. (2005b, c), Garg et al. (2023a). The Gale-substitutes property may be useful to design simple such mechanisms and combinatorial algorithms for solving the convex program (EG). For the linear case, Devanur et al. (2008) gave a primal-dual algorithm that was followed by a series of works on combinatorial algorithms for various market settings; see, e.g., Chaudhury et al. (2022), Cole and Gkatzelis (2015), Duan and Mehlhorn (2015), Garg and Végh (2019).
Appendix A Examples

### v3

@@ Appendix D Generalized network utilities
We claim that $\delta^{\prime}>1$. Recall that fitting pairs are required to satisfy (10). The sign of $\gamma_{e}\pi^{\prime}_{w}-\pi^{\prime}_{v}$ is the same as the sign of $\gamma_{e}\pi_{w}-\pi_{v}$ for all arcs $e=(v,w)$ with both endpoints inside $T$ or outside $T$. There is no arc $e=(v,w)$ entering or leaving $T$ with $0<f^{\prime}_{e}<u_{e}$, since $e,\overleftarrow{e}\in E^{\textrm{aux}}_{\mathbf{f}^{\prime},\bm{\pi}}$ for all such arcs. For every arc $e=(v,w)$ entering $T$ with $f^{\prime}_{e}=0$, we must have $\gamma_{e}\pi_{w}-\pi_{v}<0$ so that $e\notin E^{\textrm{aux}}_{\mathbf{f}^{\prime},\bm{\pi}}$. Similarly, for every arc $e=(v,w)$ leaving $T$ with $f^{\prime}_{e}=u_{e}$, we must have $\gamma_{e}\pi_{w}-\pi_{v}>0$. Hence, we can always choose $\delta^{\prime}>1$. If $\delta^{\prime}=\alpha^{\prime}$, then $\bm{\pi}^{\prime}$, $\mathbf{x}^{\prime}$, $\mathbf{q}$ and $b^{\prime}$ satisfy (11), since the right hand side for all $j\in\operatorname{supp}(\mathbf{x}^{\prime})$ has increased by a factor $\alpha^{\prime}$, and therefore equality holds for all these indices. We can therefore conclude $\mathbf{x}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b)$ and terminate by outputting $\mathbf{x}^{\prime}$.
@@
In case $\delta^{\prime}<\alpha^{\prime}$, then (11) holds with strict inequality for all $j$ for $\bm{\pi}^{\prime}$, $\mathbf{x}^{\prime}$, $\mathbf{q}$, and $b^{\prime}$. Moreover, the ratio of the right hand side and the left hand side is the same for all $j\in\operatorname{supp}(\mathbf{x}^{\prime})\cap\operatorname{supp}(\mathbf{q})$, namely, $\alpha^{\prime}/\delta^{\prime}<\alpha$. We now restart the above procedure with $(\mathbf{f}^{\prime},\bm{\pi}^{\prime})$ and $\mathbf{x}^{\prime}$ in place of $(\mathbf{f},\bm{\pi})$ and $\mathbf{x}$.
@@
We claim that the algorithm terminates in a finite number of iterations. This holds because the linear program Pullback-Flow only depends on the sign pattern of $\gamma_{e}\pi_{w}-\pi_{v}$ for the arcs $e=(w,v)$ and on $\operatorname{supp}(\mathbf{x})$. There are $3^{|E|}$ possible sign patterns; the set $\operatorname{supp}(\mathbf{x})$ is monotone decreasing and therefore there are at most $|{\mathcal{M}}|$ different possibilities. The same configuration of arc sign patterns and $\operatorname{supp}(\mathbf{x})$ cannot occur twice, because $u(\mathbf{x}^{\prime})=\operatorname{net}_{\mathbf{f}^{\prime}}(t)$ is strictly decreasing in every iteration of the algorithm. Hence, after a finite number of iterations we arrive at $\bm{\pi}^{\prime}$ and $\mathbf{x}^{\prime}$ satisfying (11) for $b^{\prime}$. Since $\mathbf{x}^{\prime}$ is decreasing throughout, we have $\mathbf{x}^{\prime}\leq\mathbf{x}$ for the input supply vector $\mathbf{x}$, as required. ∎
Input: $\mathbf{q}\in\operatorname{pdom}(u)$, $b,b^{\prime}>0$, $\mathbf{x}\in\mathcal{GD}^{u}(\mathbf{q},b)$.
Output: $\mathbf{0}\leq\mathbf{x}^{\prime}\leq\mathbf{x}$ such that $\mathbf{x}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q},b^{\prime})$.
@@
1 Find a fitting pair $(\mathbf{f},\bm{\pi})$ with respect to $\mathbf{x}$ such that $u(\mathbf{x})=\operatorname{net}_{\mathbf{f}}(t)$ and (11) holds;
2 if $u(\mathbf{x})=0$ or $\mathbf{q}=\mathbf{0}$ then return $\mathbf{x}$;
3 $\bar{\mathbf{x}}\leftarrow\mathbf{x}$ ; $\alpha\leftarrow b/b^{\prime}$ ;
@@
5 $(\mathbf{x}^{\prime},\mathbf{f}^{\prime})\leftarrow\textnormal{{Pullback-Flow}}(\bar{\mathbf{x}},\mathbf{f},\bm{\pi})$ ;
6 $\alpha^{\prime}\leftarrow\alpha\frac{u(\mathbf{x}^{\prime})}{u(\mathbf{x})}$ ;
@@ 8    $\lambda\leftarrow\frac{\alpha-1}{\alpha-\alpha^{\prime}}$ ;
9    $\mathbf{x}^{\prime\prime}\leftarrow\frac{\alpha-1}{\alpha-\alpha^{\prime}}\mathbf{x}^{\prime}+\frac{1-\alpha^{\prime}}{\alpha-\alpha^{\prime}}\bar{\mathbf{x}}$ ;
10    return $\mathbf{x}^{\prime\prime}$
@@
12    $T\leftarrow\left\{v\in V\,:\,v\text{ can reach }\operatorname{supp}(\mathbf{x}^{\prime})\text{ in }(G,E^{\textrm{aux}}_{\mathbf{f}^{\prime},\bm{\pi}})\right\}$  ;
13    $\delta^{\prime}\leftarrow\min\left\{\alpha,\min\left\{\frac{\gamma_{e}\pi_{w}}{\pi_{v}}\,:\,e=(v,w)\in\partial^{\mathrm{in}}(T)\right\},\min\left\{\frac{\pi_{v}}{\gamma_{e}\pi_{w}}\,:\,e=(v,w)\in\partial^{\mathrm{out}}(T)\right\}\right\}$;
14    if $\delta^{\prime}=\alpha^{\prime}$ then return $\mathbf{x}^{\prime}$;
@@
16    $\alpha\leftarrow\alpha^{\prime}/\delta^{\prime}$ ; $\bar{\mathbf{x}}\leftarrow\mathbf{x}^{\prime}$ ;
@@
Lemma D.8.
Let $u\,:\,\mathbb{R}_{\geq 0}^{\mathcal{M}}\to\mathbb{R}_{\geq 0}$ be a generalized network utility function and $\mathbf{q},\mathbf{q}^{\prime}\in\operatorname{pdom}(u)$ such that $\mathbf{0}\leq\mathbf{q}^{\prime}\leq\mathbf{q}$, and let $\mathbf{x}\in\mathcal{GD}^{u}(\mathbf{q},b)$. Then, there exists $\mathbf{x}^{\prime}\in\mathcal{GD}^{u}(\mathbf{q}^{\prime},b)$ such that $x^{\prime}_{j}\leq x_{j}$ whenever $q^{\prime}_{j}=q_{j}$.
@@
Clearly, it suffices to show that the statement holds when $q^{\prime}_{k}<q_{k}$ for a single good $k\in{\mathcal{M}}$. Again, we start from fitting pair $(\mathbf{f},\bm{\pi})$ with respect to $\mathbf{x}$ as in Lemma D.6. Similarly to Lemma D.7, we may assume that $u(\mathbf{x})>0$ and $\mathbf{q}\neq\mathbf{0}$. Let us now replace $\mathbf{q}$ by $\mathbf{q}^{\prime}$. The condition (11) can only be violated for $j=k$. If it is not violated, then we may return $\mathbf{x}^{\prime}=\mathbf{x}$; this is always the case when $\pi_{k}=0$. For the rest of the proof, assume that the (11) is violated for $k$ with the cost $q^{\prime}_{k}$. Note that $\pi_{j}=\beta q^{\prime}_{j}$ for every $j\neq k$, where $\beta\stackrel{{\scriptstyle\scriptscriptstyle{\mathrm{def}}}}{{=}}u(\mathbf{x})/b$; we now have $\pi_{k}>\beta q^{\prime}_{k}\geq 0$. This value of $\beta$ will be fixed throughout the argument.
We first consider the case $q^{\prime}_{k}>0$. The proof uses the subroutine $\textnormal{{Reroute-Flow}}(\mathbf{x},\mathbf{f},\bm{\pi})$ with the special node $k$. This subroutine finds $\mathbf{x}^{\prime}$ and $\mathbf{f}^{\prime}$ such that $(\mathbf{f}^{\prime},\bm{\pi})$ is a fitting pair with respect to $\mathbf{x}^{\prime}$, and further $x^{\prime}_{k}\geq x_{k}$, $x^{\prime}_{j}\leq x_{j}$ for $j\neq k$, and $u(\mathbf{x}^{\prime})=\operatorname{net}_{\mathbf{f}}(t^{\prime})\geq\operatorname{net}_{\mathbf{f}}(t)=u(\mathbf{x})$. Consider the set of nodes $T$ reachable from $k$ in $E^{\textrm{aux}}_{\mathbf{f}^{\prime},\bm{\pi}}$. According to Lemma D.5, $t\notin T$ and $T\cap\operatorname{supp}(\mathbf{x}^{\prime})=\{k\}$. Similarly as Case II in the proof of Lemma D.7, we can decrease $\pi_{v}$ multiplicatively on $T$ to obtain $\bm{\pi}^{\prime}$ such that $(\mathbf{f}^{\prime},\bm{\pi}^{\prime})$ is a fitting pair with respect to $\mathbf{x}^{\prime}$. We choose the largest multiplicative decrease such that $(\mathbf{f}^{\prime},\bm{\pi}^{\prime})$ remains a fitting pair, and that $\pi_{k}^{\prime}\geq\beta q_{k}^{\prime}$ is maintained. We then repeat this procedure for $\mathbf{x}^{\prime},\mathbf{f}^{\prime}$, and $\bm{\pi}^{\prime}$ in the place of $\mathbf{x},\mathbf{f}$, and $\bm{\pi}$.
We can again argue that this algorithm terminates in a finite number of iterations. Note that $\operatorname{supp}(\mathbf{x}^{\prime})$ is monotone decreasing throughout, and $\pi^{\prime}_{j}=\beta q^{\prime}_{j}$ is maintained for all $j\in\operatorname{supp}(\mathbf{x}^{\prime})\setminus\{k\}$. For goods with $x^{\prime}_{j}=0$, $\pi^{\prime}_{j}$ may decrease, and therefore $\pi^{\prime}_{j}\leq\beta q_{j}$ for these goods. At termination, we have $\pi^{\prime}_{j}=\beta q^{\prime}_{j}$ for all $j\in\operatorname{supp}(\mathbf{x}^{\prime})$ — including $j=k$ — and $\pi^{\prime}_{j}\leq\beta q^{\prime}_{j}$ whenever $x^{\prime}_{j}=0$. Moreover, note that $x^{\prime}_{j}\leq x_{j}$ for all $j\neq k$.
@@
Consider now the case $q^{\prime}_{k}=0$. We proceed as above, but our goal is to eventually set $\pi^{\prime}_{k}=0$. As in the previous case, we repeatedly call $\textnormal{{Reroute-Flow}}(\mathbf{x},\mathbf{f},\bm{\pi})$ and then decrease $\pi^{\prime}_{v}$ multiplicatively on $T$ by the largest possible factor such that $(\mathbf{f}^{\prime},\bm{\pi}^{\prime})$ remains a fitting pair. If this largest factor is unbounded, then we can set $\pi^{\prime}_{v}=0$ for all $v\in T$. Again, the algorithm should terminate in a finite number of iterations, using also the assumption that $\mathbf{q}^{\prime}\in\operatorname{pdom}(u)$. Once $\pi^{\prime}_{k}=0$ is reached, we again have $\pi^{\prime}_{j}=\beta q^{\prime}_{j}$ for all $j\in\operatorname{supp}(\mathbf{x}^{\prime})$, including $j=k$, and $\pi^{\prime}_{j}\leq\beta q^{\prime}_{j}$ whenever $x^{\prime}_{j}=0$. The rest of the proof is the same as in the case $q^{\prime}_{k}>0$. ∎
This concludes the proof of Theorem 1.5, namely, that all generalized network utilities are Gale-substitutes. Nevertheless, they are not $\Sigma$-Gale-substitutes, and so Theorem 1.3 is only applicable in the case when the generalized network utilities are non-satiable. We conclude by showing Theorem 1.6. This is a modification of the argument in Section B; we omit some simple details.
Proof of Theorem 1.6.
In the proof of Theorem 1.3 in Section B, property ($\Sigma$) is only used in the proof of Lemma 5.2. We show that for generalized network utilities, if $\mathbf{x}\in{\mathcal{D}}^{u}(\mathbf{p},b)$ for $\mathbf{p}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}$ and further $\mathbf{x}$ minimizes $\left\langle\mathbf{p},\mathbf{x}\right\rangle$ in ${\mathcal{D}}^{u}(\mathbf{p},b)$, then there exists $\mathbf{y}\in\mathcal{GD}^{u}(\mathbf{p},b)$ such that $\mathbf{y}\leq\mathbf{x}$. From this property, we can derive Lemma B.1 for thrifty competitive equilibria; the rest of the proof remains identical to the proof of Theorem 1.3.
Consider now (6) in the proof of Lemma 5.2. If $\lambda>0$, the same argument is applicable. Consider now the case $\lambda=0$. This means that the constraint $\left\langle\mathbf{p},\mathbf{x}\right\rangle\leq b$ is not active in the program defining ${\mathcal{D}}^{u}(\mathbf{p},b)$; that is, $\max_{\mathbf{z}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}\{u(\mathbf{z})\,:\,\left\langle\mathbf{p},\mathbf{z}\right\rangle\leq b\}=\max_{\mathbf{z}\in\mathbb{R}_{\geq 0}^{\mathcal{M}}}u(\mathbf{z})=v^{\star}$. In particular, $u(\mathbf{x})=v^{\star}$. The budget minimization property means that $\mathbf{x}$ is an optimal solution to the program
@@
$\operatorname{net}_{\mathbf{f}}(t)$ $\geq v^{\star}$
$\operatorname{net}_{\mathbf{f}}(v)$ $\geq-x_{v}\quad\forall v\in S$
$\operatorname{net}_{\mathbf{f}}(v)$ $\geq 0\quad\forall v\in V\setminus(S\cup\{t\})$
$\mathbf{0}$ $\leq\mathbf{f}\leq\mathbf{h}\,.$
From here, using LP duality and Lemma D.1, we can see that there exists a supergradient $\mathbf{g}\in\partial^{\wedge}u(\mathbf{x})$ and $\mu\in\mathbb{R}_{\geq 0}$ such that
@@
If $\mu>0$, then we can use the same argument for $\lambda=1/\mu$. If $\mu=0$, given that $\mu$ acts as the dual variable of $\operatorname{net}_{\mathbf{f}}(t)\geq v^{\star}$, which is no longer active, the optimal value of (12) becomes 0 since $\mathbf{0}$ is a feasible allocation of $u(\cdot)$ according to our definition. This implies the existence of a zero-cost optimal allocation $\mathbf{x}$: $u(\mathbf{x})=v^{*}$ and $\mathbf{p}^{\top}\mathbf{x}=0$, which indicates that $\mathbf{x}$ belongs to both ${\mathcal{D}}^{u}(\mathbf{p},b)$ and $\mathcal{GD}^{u}(\mathbf{p},b)$. Thus, the result follows. ∎
Appendix E Examples